OpenStudy (anonymous):

f(x)=1.7sin(0.018x-1.5)+12.5 I need the AMPLITUDE, PERIOD, PHASE SHIFT, and VERTICAL SHIFT.

OpenStudy (anonymous):

vertical shift is 12.5 units up phase shift is 1.5 units right but I am unure about amplitude and the period. Is the period 0.018 and amplitude 1.7 ?

OpenStudy (anonymous):

unsure

OpenStudy (anonymous):

just period is incorrect, the leftover is ok

OpenStudy (anonymous):

period = \(\dfrac{2\pi}{0.018}\)

OpenStudy (anonymous):

and I have to simplify the period or leave it as it is?

OpenStudy (anonymous):

either of them is correct, but If you simplify, let pi as it is. I mean if simplified, it should be something *pi

OpenStudy (anonymous):

so I can say\[period~ = ~\frac{π}{0.009}\]?

OpenStudy (anonymous):

oh, no, how can?? 2/0.018 is not 1/0.009 hihihi redo, please

OpenStudy (anonymous):

well, I'll leave it as it is 2π/0.018

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

tnx for the help.